# Identify As A Human Does: A Pathfinder of Next-Generation Anti-Cheat Framework for First-Person Shooter Games

### v1

@@ Abstract
The gaming industry has experienced substantial growth, but cheating in online games poses a significant threat to the integrity of the gaming experience. Cheating, particularly in first-person shooter (FPS) games, can lead to substantial losses for the game industry. Existing anti-cheat solutions have limitations, such as client-side hardware constraints, security risks, server-side unreliable methods, and both-sides suffer from a lack of comprehensive real-world datasets. To address these limitations, the paper proposes Hawk, a server-side FPS anti-cheat framework for the popular game CS:GO. Hawk utilizes machine learning techniques to mimic human experts’ identification process, leverages novel multi-view features, and it is equipped with a well-defined workflow. The authors evaluate Hawk with the first large and real-world datasets containing multiple cheat types and cheating sophistication, and it exhibits promising efficiency and acceptable overheads, shorter ban times compared to the in-use anti-cheat, a significant reduction in manual labor, and the ability to capture cheaters who evaded official inspections.
1 Introduction
@@
The gaming industry has experienced substantial growth recently, with the global game market generating \$184.0 billion in 2023 and projected to reach \$218.7 billion in 2024 [47, 48]. However, cheating in online games poses a significant threat to the integrity of gaming experiences and disrupts gameplay balance. In 2023 alone, cheating led to an estimated \$29 billion in losses, with 78% of gamers deterred from playing due to cheating [32]. In particular, first-person shooter (FPS) games, which account for 20.9% of total game sales and rank second among all genres [63], are particularly targeted by cheat developers seeking monetary gains due to their competitive and multiplayer nature. The most prevalent forms of cheating involve the use of aimbots to enhance a player’s precision and wallhacks to reveal opponents’ positions, thereby compromising game integrity [4, 40, 9].
@@
On realizing the critical importance of game security [25], both academia and industry have proposed many anti-cheats solutions, applied either on the client-side [9, 58, 49] or server-side [69, 24, 70, 2, 39, 27]. However, current solutions have the following limitations.
@@
Client-side solutions face hardware constraints, bring security and privacy concerns, or incur unacceptable system overheads. Two state-of-the-art systems [9, 49] require specific Trusted Execution Environments (TEEs) like Intel SGX. This dependence on particular hardware, which was deprecated in 2021 [5, 31], limits compatibility and practicality for industrial use. Some solutions involve scanning hard drives and gaining root privileges, posing risks of unauthorized access and personal data leakage [45, 43, 66]. Vulnerabilities, such as RCE exploits [8] and CVEs [15, 16, 17, 18], demonstrate the potential for system instability and privacy breaches. The others rely on continuous surveillance or frequent patching to counter varying cheat signatures [22, 3, 35, 58], leading to increased client-side system overheads.
@@
On the other hand, the existing server-side solutions use basic features and share low recall and accuracy. For instance, some solutions [69, 24, 70, 2, 39, 27] use features like winning rates or playtime, which are insufficient for detecting complex cheats such as aimbots or wallhacks. These methods also lack robust workflows to examine false positives, especially in noise-rich, real-world datasets. Additionally, they show low recall and accuracy in the real world, with extended ban cycles taking days or weeks [33, 50]. Incidents like the 2024 APEX Legends tournament hack [59] highlight the inefficacy of existing solutions under commercial anti-cheats such as Easy Anti-Cheat (EAC) [22].
@@
Moreover, both client-side and server-side solutions suffer from a lack of comprehensive datasets, often relying on limited or synthetic data [9, 49, 24, 70, 2, 39], which may not represent real-world cheating scenarios. The simulated datasets lack varying levels of cheating sophistication (explained in Section 2.2), which could result in performance degradation or even failure in complex real-world scenarios.
@@
To address the aforementioned limitations, we propose Hawk, a server-side approach designed to analyze data transmitted to the server, thereby avoiding the inherent limitations of client-side solutions. Hawk introduces a comprehensive set of multi-view features and a robust framework to meticulously monitor players’ points of view, statistics, and behavioral consistencies. This approach includes a corresponding workflow aimed at shortening the ban cycle and considering the occurrence of false positives. Hawk achieves high detection recall and acceptable accuracy across large, real-world datasets with varying levels of cheating sophistication.
@@
The key innovation of Hawk lies in its ability to mimic how human experts identify cheaters by focusing on three main aspects: analyzing the player’s point of view, statistically evaluating the player’s performance, and assessing the player’s gaming sense and performance consistency. Based on these aspects, we design structured and temporal features that accurately represent player behaviors. Hawk reproduces the human identification process using Long Short-Term Memory (LSTM) and attention mechanisms, ensemble learning, and deep learning networks, respectively.
@@
We evaluate Hawk on a mainstream FPS game, CS:GO. To the best of our knowledge, this paper is the first server-side work using a substantial volume of real-world datasets with different cheat types and levels of sophistication, making Hawk and the datasets more convincing for industrial use. Specifically, our dataset includes 2,979 aimbots and 2,971 wallhacks, totaling 56,041 players. The scale of the dataset has increased by two orders of magnitude compared to the state-of-the-art work [9]. Hawk achieves a maximum of 84% recall and 80% accuracy in detecting both aimbots and wallhacks, outperforming the official in-use anti-cheat. It successfully identifies cheaters who had evaded previous detections and those never caught by the official systems. We created a demo website [^1] Demo website. This is an anonymous website. to showcase some escapees’ illegal actions. Hawk validates its robustness against cheat evolution (introduced in Section 6.7) in real-world scenarios and showcases acceptable overheads. We particularly discuss Hawk’s generalizability in Section 7. The contributions of our work are as follows.
@@
  • Real-world evaluations with strong performance.
  • Open-source Hawk and dataset [^2] Link. The repository and dataset will be released upon acceptance. Artifacts are uploaded for the reference. Ethics statement is shown in Section 10. for future research.
2 Background
This section introduces anti-cheat and cheating techniques, replay systems, and replay files in modern FPS games.
2.1 Server-side and Client-side Anti-Cheat
Client-side anti-cheats focus on data, process, and hardware protection [49, 58], or detection by either examining the existing cheats’ signatures [22, 3] or classifying through behavior [9]. However, attackers can bypass it with higher privileges, by sniffing the detection report through side-channel attack [23] and altering them in the transport layer, or by changing the program’s signature [35].
Server-side anti-cheats [69, 70, 2, 24, 39] emphasize data analysis and anomaly detection. Although the cheater’s inputs might be altered on the client, the cheaters must ensure their performance outperforms others to secure the win. Thus, it is easier to detect cheats on the server because it records all clients’ final-state actions that influence the gameplay. But the prior works have limitations mentioned in Section 1.
2.2 Cheat Types in FPS Games
Wallhack and aimbot are the two most prevalent types of cheats in FPS games [1]. Both types of cheats are mainly done by accessing the client’s memory data or image information. Wallhack enables players to visually penetrate obstructions like walls, which allows players to spot opponents across the map and effortlessly track their movements, providing a significant tactical advantage. Aimbot assists a player’s aims and shots so that the cheaters gain transcendent precision and reaction speed. Aimbot’s functionality can be customized according to player preference, allowing one to obtain different aiming performance, ranging from brutal-force aims to smooth aims like human [9]. Aimbot can be subdivided into pure aimbot (the primitive aimbot), triggerbot (auto-shoot only when the cross-hair is on the opponent), micro-settings (counteract recoil with Hotkeys), and computer-vision-based aimbot (aim with object detection models). In the following sections, aimbot refers to all aforementioned subcategories.
Notably, cheating provides different levels of sophistication [66]. The cheater might cheat cautiously, pretending to be a normal player, switching the cheat on and off, or cheating unabashedly with different configurations, various brands of cheating software, and categories. However, cheaters’ behaviors will always differ from the normal players [9]. Thus as long as cheaters want to gain unfair advantages, they must outperform the average. This is bound to create anomalies at the data level and therefore should be identifiable by models. Our evaluations in Section 6 include different levels of cheating sophistication by using large real-world datasets.
2.3 Replay System and Replay File
@@ 3 Overview
3.1 Fact Observations
Although the anti-cheat tasks are still challenging for automation systems in not just the FPS field but almost any game genre, the illegal activities are identifiable for seasoned human players [11]. Therefore, by harnessing such inspiration, we discovered three observations that human experts always focus on to identify cheaters. {mdframed}[style=obs] Observation 1 (OBS.1). Review a player’s point-of-view (POV). Akin to watching a player’s game replay unfold over time, this review process involves analyzing various in-game behaviors, such as aim and shot, positioning and movement, and the in-game economy management and props utilization. These behaviors can provide temporal insights into whether the player is legitimately playing the game. {mdframed}[style=obs] Observation 2 (OBS.2). Conduct an in-depth analysis of a player’s statistical data. Cheaters often exhibit statistical anomalies that set them apart from normal players. These anomalies include abnormally high accuracy and eliminations, a low death rate, and a high number of kills made through obstacles, etc. These statistical deviations serve as strong indicators of cheating. {mdframed}[style=obs] Observation 3 (OBS.3). Examine the consistency between a player’s game sense and combat performance. A player requires substantial game experience to have better performances, which results in a positive correlation between a player’s game sense and their performance. Thus, the inconsistency between the two aforementioned factors indicates potential cheating. For instance, if a player moves like a rookie but shoots and aims incredibly accurately, it is suspicious.
3.2 Hawk’s Workflow
Figure 1 illustrates Hawk’s workflow, comprising four interlinked stages.
Stage One. Preprocessing. The demo is applied to parse in-game data from the banned databases. The multi-view features are then extracted from parsed data and then annotated for training. We discuss the feature constructions in Section 4.
Stage Two. Hawk deployment. Hawk utilizes the extracted features as inputs for its subsystems RevPov (Section 5.1), RevStats (Section 5.2), and ExSPC (Section 5.3) to mimic human identification process, respectively. The outputs are incorporated into Mvin subsystem (Section 5.4) that presents a final cheating report.
Stage Three. Game Master (GM) team verification. Given the potential personal assets tied to a player’s account, a ban should not be issued solely based on an algorithmic verdict. It is imperative to provide substantial evidence before initiating such an action. Hence, the final verdict requires GMs’ verification to avoid false bans.
Stage Four. Data updates. Upon the affirmation of cheating, the relevant replay files are added to the banned databases. This act enriches and updates the reservoir of training samples which may contain novel patterns and behaviors. Thus, four stages constitute a cyclical workflow, with each stage contributing to the subsequent one.
This ongoing cyclical workflow enhances the system’s robustness and provides a dynamic solution to multi-type cheat detection. We further discuss this in Section 6.7
@@ 3.2 Hawk’s Workflow
4 Feature Constructions
The feature design is one of our contributions and key to the server-side anti-cheat. This section introduces the construction of temporal features, structured features, and the classification of sense and performance features. These features are intended to be a comprehensive representation of the three aspects of human identification of cheating in Section 3.1.
4.1 Temporal Features
Temporal features are information derived from the demo in the time domain. Each temporal data tuple represents the state information (e.g., coordinates, view directions, events, etc.) associated with the present tick. Where tick denotes the smallest unit of time in the game. This section describes temporal features’ components and illustrates the distinctions in temporal features between normal players and cheaters.
4.1.1 Temporal Features Construction
According to different in-game behaviors, temporal features are segmented into three main categories and subdivided into seven specific types. This overview provides only categories and types due to the less-expressive nature of temporal features. Appendix B lists elaborate descriptions.

### v2

@@ Abstract
The gaming industry has experienced substantial growth, but cheating in online games poses a significant threat to the integrity of the gaming experience. Cheating, particularly in first-person shooter (FPS) games, can lead to substantial losses for the game industry. Existing anti-cheat solutions have limitations, such as client-side hardware constraints, security risks, server-side unreliable methods, and both-sides suffer from a lack of comprehensive real-world datasets. To address these limitations, the paper proposes Hawk, a server-side FPS anti-cheat framework for the popular game CS:GO. Hawk utilizes machine learning techniques to mimic human experts’ identification process, leverages novel multi-view features, and is equipped with a well-defined workflow. Hawk is evaluated with the first large and real-world datasets containing multiple cheat types and cheating sophistication, and it exhibits promising efficiency and acceptable overheads, shorter ban times, higher recall and similar false positive rate compared to the in-use anti-cheat, and the ability to capture cheaters who evaded official inspections.
I Introduction
@@
The gaming industry has experienced substantial growth recently, with the global games market generating \$187.7 billion in 2024 and projected to generate \$213.3 billion by 2027 [1]. However, cheating in online games poses a significant threat to the integrity of gaming experiences and disrupts gameplay balance. In 2023 alone, cheating led to an estimated \$29 billion losses, with 78% of players discouraged from playing due to cheating [2]. In particular, first-person shooter (FPS) games, which account for 20.9% of total game sales [3], are particularly targeted by cheat developers because their stringent latency requirements make key computations on the client side, creating fertile ground for cheats to exploit.
The most prevalent forms of cheating involve the use of aimbots to aid aiming and wallhacks to reveal opponents’ positions, thus compromising game integrity [4, 5, 6].
@@
Realizing the critical importance of game security [7], both academia and industry have proposed many anti-cheats solutions, applied either on the client side [6, 8, 9] or server side [10, 11, 12, 13]. However, current solutions have the following limitations.
@@
Client-side solutions face hardware constraints, security and privacy concerns, or unacceptable system overheads. Two state-of-the-art systems [6, 9] require a specific Trusted Execution Environment (TEE), Intel SGX. This dependence on particular hardware, which was deprecated in 2021 [14, 15], limits the compatibility and practicality for industrial use. Some industrial solutions [16, 17, 18] involve scanning hard drives and gaining root privileges, posing risks of unauthorized access and personal data leakage [19, 20, 21]. Vulnerabilities, e.g., RCE exploits [22] and CVEs [23, 24, 25, 26], demonstrate the potential for system instability and privacy breaches. And most client-side works rely on concurrent surveillance [9, 8] leading to increased client-side overheads, or frequent patching to counter varying cheat signatures [16, 17, 18, 6], which requires extra engineering efforts and is considered to be less effective [13] since it only identifies known and deprecated cheats. Importantly, client-side data and decision results are unreliable and easily tampered with [27, 28, 29, 13]. Thus, the industry has shifted towards server-side anti-cheat solutions [27, 28].
@@
However, existing server-side solutions use basic features and have low recall and accuracy. For instance, previous works [10, 11, 13, 30, 31] use naive features like winning rates, playtime, or count shots on an invisible bait which are insufficient or inapplicable for detecting today’s complex cheats or specific cheat types. These works leverage thresholds to classify, which is not robust enough for practical use, especially in noise-rich, real-world datasets. Additionally, some commercial anti-cheats integrate such server-side approaches [16]. Yet, they show low recall and accuracy, with extended ban cycles and increasing community grievances [32, 33]. Incidents like the 2024 APEX Legends Tournament hack [34] and the study [30] highlight the ineffectiveness of existing commercial anti-cheats, e.g., EAC [16] and VAC [18].
@@
More broadly, prior research on anti-cheat techniques suffers from the lack of comprehensive real-world datasets, with most studies relying on limited or synthetic data [6, 9, 10, 11, 12, 13], which are hard to capture diverse cheating scenarios and varying levels of sophistication (explained in Section II-B). This could result in performance degradation or even failure in complex real-world scenarios.
@@
To address the aforementioned limitations, we propose Hawk, a server-side approach designed to analyze data transmitted to the server, thereby avoiding the inherent limitations of client-side solutions. Hawk introduces a comprehensive set of multi-view features and a robust framework to meticulously monitor players’ points of view, statistics, and behavioral consistencies. This approach includes a corresponding workflow that aims to shorten the ban cycle and considers the occurrence of false positives. Hawk achieves high detection recall and acceptable accuracy in large real-world data sets with varying levels of sophistication of cheating. Table I listed the comparison between Hawk and other detection methods.
TABLE I: Detection works comparison. ‘Availability’: public availability of technical details (T), source code (C), and datasets (D), red and green indicate unavailable and available. ‘Dynamic Tuning’: ability to adjust different anti-cheat demands (e.g., prioritize different metrics, introduced in Section V-D) without retraining the entire system. ‘Data Extraction-Game Coupling Degree’: difficulty of acquiring the game’s raw data, ‘high coupling’ indicates the process is highly coupled with the game logic (e.g., modify engine, hooking process), making it challenging to generalize or reuse across different game versions.
Side Method Work(s) Availability
Detection Feature
Comprehensiveness
Data Extraction-Game
Coupling Degree
Real-world Functionality
Dataset Deployment Constraint Robustness Dynamic Tuning Performance
Client ML-based [6] T, C, D Single aspect & data type High Coupling SSL1 TEE (Intel SGX) Unknown Unknown2
Commercial signature-based [16, 17, 18, 35, 36] T, C, D Unknown High Coupling Real-world OS and kernel drivers Unknown Unknown Low recall, acceptable FPR
Server Statistical/Threshold-based [10, 11, 13, 30, 31] T, C, D Single aspect & data type High Coupling SSL Concurrent overheads Unknown Unknown
ML-based [12] T, C, D Single aspect & data type High Coupling SSL Concurrent overheads Unknown Unknown
Hawk T, C, D Multiple aspects & data types Low Coupling Real-world None CE3, N4 High recall, similar FPR
1SSL: Simulated, small-scale, low complexity; 2Unknown: Never been tested or disclosed; 3CE: Robust to cheat evolution (an adversarial attack in Section VI-G); 4N: Robust to noise.
@@
The key innovation of Hawk lies in its ability to mimic how human experts identify cheaters by focusing on three dimensions: analyzing the player’s point of view, statistically evaluating the player’s performance, and assessing the player’s gaming sense and performance consistency. Building on these dimensions, we design structured and temporal features that precisely describe player behaviors. These features are selected from hundreds of candidates to ensure that they effectively represent diverse behavioral patterns. Hawk replicates the human identification process by leveraging machine learning techniques, including Long Short-Term Memory (LSTM) with attention mechanisms, ensemble learning, and deep neural networks, each tailored to specific aspects of the analysis. Hawk addresses the challenges of integrating heterogeneous data structures and dynamically adjusting anti-cheating priorities by developing a dedicated integration subsystem, Mvin.
@@
We evaluate Hawk on a mainstream FPS game, CS:GO. To the best of our knowledge, this paper is the first server-side work using a substantial volume of real-world datasets with different cheat types and levels of sophistication, making Hawk and the datasets more convincing for industrial use. Specifically, our dataset includes 2,979 aimbots and 2,971 wallhacks, totaling 56,041 players. The scale of the dataset has increased by two orders of magnitude compared to the state-of-the-art work [6]. Hawk achieves at most 84% recall and 80% accuracy in detecting aimbots and wallhacks. Hawk outperforms baselines in the separate comparisons across different subsystems, given Hawk’s diverse input data structures. It successfully identifies cheaters who had evaded previous detections and those who were never caught by official systems. We created a demo website [^1] Demo website: https://hawk-anticheat.github.io/. to showcase some escapees’ illegal actions. Hawk validates its robustness against cheat evolution (introduced in Section VI-G) in real-world scenarios and showcases acceptable overheads. We particularly discuss Hawk’s generalizability in Section VII. The contributions of our work are summarized below.
@@
  • Real-world data evaluations with strong performance.
  • Open-source Hawk and dataset [^2] Public repository and dataset: 10.6084/m9.figshare.25940818. Ethics considerations are shown in Supplemental Material A. for future research.
II Background
The section introduces anti-cheat, cheating techniques, the current state of cheating, and the replay design in FPS games.
II-A Server-Side and Client-Side Anti-Cheat
Client-side anti-cheats focus on data, process, and hardware protection [9, 8], or detection by examining the existing cheats’ signatures [16, 17] or classifying through aiming behavior [6]. Yet, attackers can bypass it with higher privileges, by sniffing the detection report through side-channel attack [37] and altering them in the transport layer, or by changing the program’s signature [38]. Server-side anti-cheats [10, 11, 12, 39, 13, 31] emphasize data analysis and anomaly detection. Although the data might be altered on the client, they must secure victory by outperforming other players. Thus, detecting cheats on the server is easier because it records all clients’ final-state actions influencing the gameplay. But the prior works have the limitations mentioned in Section I.
II-B Cheating in FPS Games
Wallhack and aimbot are the two most prevalent types of cheats in FPS games [40]. Both types of cheats are mainly done by accessing the client’s memory data or image information. Wallhack enables players to visually penetrate obstructions like walls, which allows players to spot opponents across the map and effortlessly track their movements, providing a significant tactical advantage. Aimbot assists a player’s aims and shots so that the cheaters gain transcendent precision and reaction speed. Aimbot’s functionality can be customized according to player preference, allowing one to obtain different aiming performance, ranging from brutal-force aims to smooth aims like human [6]. Aimbot can be subdivided into pure aimbot (the primitive aimbot), triggerbot (auto-shoot only when the cross-hair is on the opponent), micro-settings (counteract recoil with Hotkeys), and computer-vision-based aimbot (aim with object detection models). In our paper, aimbot refers to all aforementioned subcategories. Cheating provides different levels of sophistication [21]. The cheater might cheat cautiously, pretending to be a normal player, switching the cheat on and off, or cheating unabashedly with different configurations, various brands of cheating software, and categories. However, cheaters’ behaviors will always differ from the normal players [6]. Thus, cheaters must outperform the average to gain unfair advantages. This is bound to create anomalies at the data level and therefore should be identifiable by models. Our evaluations in Section VI include different levels of cheating sophistication.
II-C Replay System and Replay File
@@ III Overview
III-A Fact Observations
Although the anti-cheat tasks are still challenging for automation systems in FPS games, the illegal activities are identifiable for seasoned human players [41]. Therefore, by harnessing such inspiration, we discovered three observations that human experts always focus on to identify cheaters. Observation 1 (OBS.1). Review a player’s point-of-view (POV). Akin to watching a player’s game replay unfold over time, this review process involves analyzing various in-game behaviors, such as aim and shot, positioning and movement, and the in-game economy management and props utilization. These behaviors can provide temporal insights into the legitimacy of a player’s gameplay. Observation 2 (OBS.2). Conduct an in-depth analysis of a player’s statistical data. Cheaters often exhibit statistical anomalies that set them apart from normal players. These anomalies include abnormally high accuracy and eliminations, a low death rate, and a high number of kills made through obstacles, etc. These statistical deviations serve as strong indicators of cheating. Observation 3 (OBS.3). Examine the consistency between a player’s game sense and combat performance. A player requires substantial game experience to have better performances, which results in a positive correlation between a player’s game sense and their performance. Thus, the inconsistency between the two aforementioned factors indicates potential cheating. For instance, if a player moves like a rookie but shoots and aims incredibly accurately, it is suspicious.
III-B Workflow
Figure 1 illustrates Hawk’s interlinked workflow, which enhances the system’s robustness and provides a dynamic solution to multi-type cheat detection (discussed in Section VI-G).
[…]
@@ III-B Workflow
IV Hawk’s Feature Design
The feature design is one of our contributions. This section introduces the construction of structured features, temporal features, and the classification of sense and performance features. These features are selected among hundreds of our designed naive features and intended to be comprehensive representations of the three aspects of human identification of cheating in Section III-A to overcome the challenge of the potentially biased classifications brought by the monospecific features in the prior works mentioned in Section I.
IV-A Structured Features
Structured features are single-value data calculated with well-designed algorithms to represent integrated in-game behavioral information per match and ensure time complexity. This section describes the components of structured features and visualizes the structured feature differences between normal players and cheaters.
IV-A1 Structured Features Construction
The structured features for Hawk are categorized into four main categories that comprise 28 features. For the sake of brevity and coherence, only the introduction of categories is outlined below. A meticulous breakdown is shown in Supplemental Material C.
  • Elimination features emphasizes metrics related to in-game kill events. For example, the occluder-penetration index is a weighted index that reflects the extent of the average occluder-penetration eliminations (e.g., eliminate the opponent behind walls or smoke) conducted by the player. In terms of the cheaters exploiting wallhacks, this index is more significant than the normal players’.
Figure 2: Box plots of comparative visualization between honest players and cheaters on top-12 structured features under the Mann-Whitney U test with distinction descending order.
IV-A2 Structured Data Comparative Visualization
Figure 2 offers a box-plot on structured features to analyze between honest players and cheaters comparatively. Each feature on the x-axis (we select top-12 features for illustration purposes) is ordered based on the outcomes of the Mann-Whitney U test, with the leftmost features, such as akpr (introduced in Supplemental Material C), marking pronounced disparities between the two groups. This implies that features towards the left end, showcased by their lower p-values, are more distinguishable, whereas those on the right are less so. The y-axis is normalized values between 0 and 1. The box plots, distinguished by light blue for honest players and hatched red for cheaters, provide insights into the median, interquartile range, and outlier distributions. In which, diamond shapes represent outliers, emphasizing the spread and distribution of data across the features. Despite considering dimensionality reduction techniques, such as PCA (Supplemental Material Figure 4), empirical tests show an optimal performance when all 28 features were employed. Therefore, cheaters often exhibit deviations in their static data compared to normal players. Even when cheaters deliberately attempt to conceal their behavior, the subtle differences tend to accumulate over time and reveal disparities statistically.
IV-B Temporal Features
Temporal features are derived in the time domain. Each temporal data tuple is the state information (e.g., coordinates, view directions, events, etc.) associated with the present tick. Where tick denotes the smallest unit of time in the game.
IV-B1 Temporal Features Construction
According to different in-game behaviors, temporal features are segmented into three main categories and subdivided into seven specific types. Supplemental Material D lists elaborate descriptions.